OpenStudy (anonymous):

A bag contains 2 gold coins and 5 silver coins. A number cube has the numbers 1 through 6 on its faces. What is the probability of drawing a silver coin from the bag and rolling an even number on the number cube?

OpenStudy (anonymous):

@hartnn

OpenStudy (***[isuru]***):

the probability of getting a silver coin is 5/7 ( since there r 5 silver coins) the probability of getting an even number on a dice is 3/ 6 ( since there r 3 even numbers on a dice) so the probability of getting those both is (5/7)x(3/6) = 15/42= 5/14

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

\[P(Silver AND Even) = P(Silver) * P(Even)\] \[=\frac{5}{7} * \frac{ 3 }{ 6 }\]\[=\frac{15}{42}\]\[=\frac{5}{14}\]